the average price of a home in tokyo is $ 335,000. in freiburg, $ 135,000. in both cities, the average price of a home is six times the average income. about how much more is the average income in tokyo than the average income in freigburg?
Answer by stanbon(75887)

the average price of a home in tokyo is $ 335,000. in freiburg, $ 135,000. in both cities, the average price of a home is six times the average income. about how much more is the average income in tokyo than the average income in freigburg?
---------------
Average Tokyo income: 335,000/6 = 55833.33
Average Frieburg income: 135000/6 = 22500
---------------------
Difference 55833-22500 = $33,333,00
=========================================
Cheers,
Stan H.